OpenStudy (anonymous):

can anyone tell me the formula for a Triangular Prism, a cone, a Rectangular Prism, and a pyramid?

OpenStudy (anonymous):

so do you mean volume or surface area?

OpenStudy (anonymous):

surface area

OpenStudy (anonymous):

ok so surface area for rectangular prism: SA=2(wl+hl+hw) triangular prism: SA = wh + lw + lh + ls cone: SA = πr2 + πrl so im assuming square based pyramid: SA = 2bs+b^2

OpenStudy (anonymous):

thanks so much!!